OpenStudy (anonymous):

The solubility of AgCl(s) in water at 25 C is 1.33 x 10^-5 mol/L and its deltaHo of solution is 65.7 kJ/mol. What is the solubility at 50.0 C?

OpenStudy (aaronq):

find Ksp, then use van't hoff equation to find Ksp at other temperature, then use the Ksp equation to find the molar solubility \[\ln(\frac{ K _{2} }{ K _{1} })=\frac{ -\Delta H }{ R }(\frac{ 1 }{ T _{2} }-\frac{ 1 }{ T _{1} })\]